Q23) What voltage applied across a 100 Ω resistor would result in a 0.25 W power dissipation? For this auto-graded question, use at least 2 significant figures with your answer in units of Volts, not mV. Also do not include units in the answer box.  

Put answer here:
Answer #1

Given Resistor value is 100 ohm

power dissipation in resistor is 0.25 W

Power dissipation in resistor is P = V^2 / R

0.25 = V^2 / 100

0.25 * 100 = V^2

25 = V^2

V = sqrt(25)

V = 5.00